### v1

# Explainable Deep Learning Analysis for Raga Identification in Indian Art Music

Abstract
The task of Raga Identification is a very popular research problem in Music Information Retrieval. Few studies that have explored this task employed various approaches, such as signal processing, Machine Learning (ML) methods, and more recently Deep Learning (DL) based methods. However, a key question remains unanswered in all of these works: do these ML/DL methods learn and interpret Ragas in a manner similar to human experts? Besides, a significant roadblock in this research is the unavailability of ample supply of rich, labeled datasets, which drives these ML/DL based methods. In this paper, we introduce ”Prasarbharti Indian Music” version-1 (PIM-v1), a novel dataset comprising of 191 hours of meticulously labeled Hindustani Classical Music (HCM) recordings, which is the largest labeled dataset for HCM recordings to the best of our knowledge. Our approach involves conducting ablation studies to find the benchmark classification model for Automatic Raga Identification (ARI) using PIM-v1 dataset. We achieve a chunk-wise f1-score of 0.89 for a subset of 12 Raga classes. Subsequently, we employ model explainability techniques to evaluate the classifier’s predictions, aiming to ascertain whether they align with human understanding of Ragas or are driven by arbitrary patterns. We validate the correctness of model’s predictions by comparing the explanations given by two ExAI models with human expert annotations. Following this, we analyze explanations for individual test examples to understand the role of regions highlighted by explanations in correct or incorrect predictions made by the model.
Index Terms:
Hindustani Classical Music, Raga Identification, Deep Learning, Explainable AI, Music Information Retrieval
I Introduction
Indian Art music (IAM) is a culturally rich and complex musical tradition that has evolved over centuries. It is categorized into two main categories: Hindustani Classical Music (HCM) and Carnatic music (CM). HCM is predominant in the northern regions of India, while Carnatic music is more prevalent in the south. Both forms of music have their own distinct nature and way of singing. Central to IAM is the concept of a Raga, which can be termed as melodic structures that provide a framework for improvisation and composition, making it fundamental descriptor for IAM. It represents the mood of any musical composition. Ragas also form the basis of a lot of indian film or folk music[1]. Detailed description about IAM Ragas can be found at [2, 3].
To compare with western music, in HCM, a raga is not merely defined by its set of notes, but also by the way these notes are performed, with emphasis on smooth melodic glides and the space between notes. This contrasts with Eurocentric music genres, where a scale is primarily defined by its set of notes[4]. Also, in western music, there is more emphasis on harmony, whereas HCM focuses on melody[5]. Raga is also different from a mode in Eurocentric music as it encompasses a broader range of musical features including scale, notes’ sequence, hierarchy, relative duration, micro-tonal intonation, ornamentation, melodic movement, specific phrases, and mood[4]. Due to these distinctions, the task of Automatic ARI in IAM differs significantly from genre classification in Western music. In IAM, categories such as Dhrupad, Khayal, Thumri, Dadra, Tappa, and Bhajan etc can be considered distinct genres[5], but not Ragas.
IAM’s emphasis on improvisation within the framework of Raga underscores the intricate nature of this musical tradition, setting it apart from Western music genres.
For data-driven methods used in computational musicology and music information retrieval (MIR), digital music artifacts with metadata and expert annotations are a prerequisite[6]. Today, searching for music based on identifiers like ragas is not as simple as searching for text on search engines like Google. Efforts are underway to automate this process, starting with the automatic identification of ragas. Music enthusiasts and researchers who view music as more than just entertainment require a music/audio retrieval system based on melody or raga[7]. ARI has several other applications, including music recommendation, music teaching, and automatic annotation.
The progress of research in IAM is often hindered by the scarcity of a variety of open labeled datasets. Introduction of new datasets in IAM can be useful for making tasks like melody extraction, ARI, tala identification, source separation, automatic mixing, and performance analysis more robust and generalisable.
For the task of ARI, various Machine Learning(ML)/deep learning(DL) based approaches are being used, but we have no way to know if the models are making predictions as if an IAM expert would do or not. This motivates us to look for these answers from the field of Explainable AI (ExAI), which has barely been explored in the field of IAM analysis. ExAI is a popular research area within ML that aims to provide explanations about complex deep learning models, making them interpretable. ExAI focuses on explaining the behavior of models and their predictions, in various known and unknown situations, in a post-hoc manner, hence increasing transparency and understanding of these models, ultimately improving their reliability and trustworthiness. In this study, we take up the task of ARI for IAM and then use ExAI methods to show the reliability of the classifier.
The Key contributions of this work are as follows:
  • We Introduce a novel HCM dataset along with manual annotations, and benchmark the dataset for the task of Raga classification by undergoing Ablation Studies. We release this dataset in the form of 30-second audio chunks along with labels and additional metadata. The sample audios can be accesed at: https://github.com/ParampreetSingh97/PIM_v1_ExAI.git.
  • We make one of the first attempts on using ExAI techniques for raga identification, and subsequently validating the significance and correctness of the Raga classifier musically.
  • We show that the ARI model in fact does make musically correct predictions, making it usable for various HCM tasks like music teaching, automatic labeling, music suggestion and search, and other tasks involving IAM Ragas.
II Related Work
II-A HCM Datasets
The Indian Art Music Raga Recognition Dataset (RRD)[8], offers 116 hours of HCM with 300 recordings featuring 30 unique ragas, with 10 recordings per raga. Then there is Saraga dataset [9], which comprises of 108 recordings of HCM in 61 unique Ragas and 9 unique talas, totaling 43.6 hours. Both of these datasets feature live performances by various experts, along with providing labels like Raga names, manually marked melodic phrases, and features like tonic and pitch extracted using various algorithms. While Saraga allows download access with annotations, RRD requires access requests for audio file viewing. Notably, we observe a significant overlap of audio files between Saraga and RRD. Some other works [10, 11] utilize self-made HCM datasets for ARI, but these datasets have not been made publicly available.
Our dataset stands out as the largest labeled HCM dataset to the best of our knowledge, surpassing both Saraga and RRD, in terms of total duration, number of recordings, unique Ragas, talas and manually labeled Raga and tonic information.
II-B Raga Identification
For HCM, the paper [10] uses their own generated instrumental dataset of 20 second snippets for 12 Raga classification. They make use of a combination of features like Pitch, MFCC and other statistics derived from mfcc together for the classification task. Another paper [4] does 2-Raga classification using a subset from Saraga Dataset and report that using Swargram features is helpful for Raga classification, which is nothing but Chromagram features shifted in pitch. The paper [12] uses pitch class distribution and N-gram approach on the extracted pitch for raga identification and reports that longer audio clip length is beneficial for the task. Another work [13] uses a dataset downloaded from youtube for the task of 6-class classification. In paper [14], KNN and SVM machine learning algorithms were used for classifying different ragas like Yaman and Bhairavi, and the KNN was found to be performing slightly better. The paper [15] proposed a hierarchical deep learning system called PhonoNet, which used a CNN model in the first stage to predict the raga of a small audio chunk, and a two-stage recurrent LSTM network in the second stage to look at various parts of the audio input and prioritize observed patterns.
There are some other works related to Carnatic Music[1, 16, 17, 18, 19, 20, 21, 22], and most of them use basic features like extracted pitch, MFCCs, Spectral centroid, chromagram, spectrogram, scalogram etc and apply different ML classifiers like Naive Bayes, SVM, Random forest, GMMs, and some also use DL models like RNN, LSTM, CNN etc.
We observe that we have more research centered around Carnatic music, primarily because of higher availability of labeled datasets as compared to HCM. Also, to the best of our knowledge, none of the works in the literature have explored ExAI methods for the task of Raga Identification in any form.
II-C ExAI methods
Until now, we have a lot of research already available focused on ExAI methods for DL models used for various tasks involving Images, and some works for audio and music also. We will now discuss some notable works in both. One very prominent kind of approach for post-hoc explainability involves the use of model’s Gradients from various layers of a CNN or ANN model and using them in a intelligent way to visualize explanations of the classifier in the form of saliency map, superimposed over the original image itself, highlighting the most important region for classification by the classifier [23, 24, 25, 26, 27, 28, 29, 30]. We will be using GradCAM++[28], which is a very popular method for CNN explanations. It weighs each of the pixels in each feature map of the last CNN layer separately, with weights determined using the gradients of these layers in a clever way, and then finally stacking all the weighted feature maps together to get the final attribution map, which highlights important regions in an image. In the audio domain, we have some works  [31, 32] which treat audio feature representations as images and use the ExAI methods for images to generate explanations.
Other type is of black box models, whose gradients are not accessible and we have to design ExAI methods by querying the model multiple times. There are some notable works for black box model explainability, with LIME[33] generating locally faithful explanations by training an interpretable model on perturbed instances of the input data sample. Another method SHAP[34], uses game theory to assign a unique contribution score to each feature in the input, indicating its impact on the model’s output. SHAP is computationally intensive, especially for models with a large number of features. Another approach Anchors[35] express explanations as simple IF-THEN rules, making them more easily understandable, but it struggles with large and complex datasets.
For audio , we have a variant of LIME in SoundLIME[36], which works on spectrograms or other audio based features, treats them as image data only and then replicate the LIME framework.
For our work, we employ a very popular white-box ExAI model in Grad-CAM++[28], and we use SoundLIME out of the black box models. We employ both white-box and black-box ExAI models to ensure comprehensive coverage of interpretability techniques, enabling a thorough comparison and analysis of the performance of our ARI classifier model.
III Dataset
The dataset comprises a total of 501 HCM songs with each song being a polyphonic audio containing vocals of the main artist, percussion instruments like tabla, Pakhawaj, along with Manjeera in some audios, along with non-percussion instruments like Harmonium, Sarangi, Violin, and also tanpura (drone) which provides the tonic for the whole music performance. These music recordings are sung in variety of Ragas and Talas with all of the songs being of varying lengths. The dataset includes 144 unique Ragas performed by 169 vocal artists. The Distribution of Ragas is shown in table: I. Harmonium is the most common instrument, present in 372 songs, while some performances feature multiple string instruments. All of the songs include drone (tanpura) playing in the background of the performance. The dataset includes performances in 20 different talas accompnying the vocals, with teentaal (121 instances) and ektaal (146 instances) being the most frequently used talas as shown in fig:1. Each song is available in the form of 30-second chunks, totaling 23,005 audio files, approximately 191.70 hours of music.
Figure 1: Distribution of number of audio files per tala
III-A Musical Styles
The dataset covers a variety of musical styles belonging to HCM, such as Khayal, Dhrupad, Dadra, Thumari, Tarana, Gat, Bhajan, Tappa, Dhammar, Kajri, and Jhoola (folk), showcasing a wide range of songs from different regions throughout India.
III-B Recording Source and Quality
All recordings are sourced from the archives of Prasar Bharti (Akashwani, Delhi), which are not available online elsewhere. They feature performances by HCM ustaads (experts) and are mainly studio recordings or recordings from live shows, ensuring high audio quality. Most of the featured vocal, percussion, or string instrument artists are well-known in the HCM community. All audio files are mono tracks. Some of the recordings feature speech at the start which includes introduction of the artists and some details about their performance and ranges from a few seconds to about a few minutes in some recordings.
III-C Metadata
The metadata contains details such as the names of vocal and instrumental artists, the instruments played, the event where the performance took place, the name of the song, and the number of 30-second files for each song. We also add to the metadata the Raga name, the tonic and the start and end times of music and speech for each of those songs, labeled manually by our annotators.
III-D Manual Annotation Process
Two music experts with over 10 years of experience in vocal music teaching and singing for HCM carried out the task of annotation. They identified the Raga and tonic of each of the 501 music performances, listed in the metadata. This annotation task involves labelling the whole audio with a single Raga and Tonic label, as they do not change throughout a single HCM performance. As the task does not involve frame-wise labelling or subjective labelling, there is negligible chance of discrepancies among annotators, so inter-annotator agreement is not necessary.
TABLE I: Distribution of number of audio files per Raga
No. of files Ragas
1 Abhogi_kanada, Alhaiya_bilawal, Asavari, Asavari (Komal_rishabh), Basant_bhairav, Bhimpalasi, Bhoop_nat, Dagori, Deepak, Des_and_tilak_kamod_combination, Deshi, Desi_todi, Devgandhar, Devgiri_bilawal, Gaud_sarang, Gunkali, Jhoola, Kafi_malhar, Kans_gandhar, Kaushiki_kanada, Lalit, Madhmad_sarang, Malgunji, Mian_ki_todi, Mishra_gara, Mishra_jhinjhoti, Mishra_jinjhoti, Mishra_kirwani, Nat_bhairav, Nat_bihag, Nat_kedar, Pancham, Prabhat_bhairav, Purvi, Ramkali, Sahna_kanada, Salagwarali, Shahana_kanada, Shiv_kalyan, Shivranjani, Shuddha_bairav, Shuddha_basant, Shuddha_sarang, Sohni (pancham), Suha_malhar, Tilak_shyam, Todi, Todi (ahiri), Vrindavani_sarang
2 Anand_bhairav, Bageshri_bahar, Bahar, Bhairav, Bhatiyar, Bihari, Chandrakauns (Bageshri_ang), Chandrakauns (Bhairavi_ang), Charukeshi, Darbari, Deshkaar, Hemant, Hindol, Jait_kalyan, Jayant_malhar, Kirwani, Marwa, Paraj, Paraj_kalingda, Pilu, Puriya_dhanashri, Raysa_kanada, Sampoorna_malkauns, Shahana, Shivmat_bhairav, Swanandi, Tilak_kamod
3 Abhogi, Bageshri_kanada, Basant_bahar, Basant_kedar, Bhinna_shadaja, Durga, Gavati, Gujri_todi, Hameer, Hansdhwani, Janasammohini, Jhinjhoti, Kafi_kanada, Kaunsi_kanada, Mishra_bhairavi, Mishra_desh, Mishra_pahadi, Mishra_tilak_kamod, Mishra_tilang, Sawani_bihag, Shyam_kalyan, Tilang
4 Aheer_bhairav, Basant, Bihagda, Gaud_malhar, Jaunpuri, Kamod, Kaushik_dhwani, Majh_khamaj, Megh, Nand, Rageshree, Saraswati, Surdasi_malhar
5 Adana, Gorakh_kalyan, Madhukauns, Mishra_pilu, Puriya
6 Chayanat, Darbari_kanada, Jaijaiwanti, Kalavati, Mian_malhar, Shankara
8 Chandrakauns, Des, Kafi, Pahadi
9 Kedar, Mishra_khamaj, Puriya_kalyan
10 Jog_kauns, Yaman
11 Malkauns, Maru_bihaag
12 Bageshri, Bhopali, Bihag, Nayaki_kanada, Shuddha_kalyan
13 Khamaj
14 Jog
15 Bhairavi, Sohni
IV Methodology
In this study, we use our PIM dataset to extract chromagram features from the audios and train a CNN-LSTM model to perform multi-class classification for 12 Raga classes. Next, we employ Grad-CAM++ (GC) and SoundLIME (SL), two popular ExAI models, for saliency visualisation and comparing them with annotations from music experts. The sub-sections below explain the entire procedure.
IV-A Feature Extraction
We use chromagram features, extracted from the audios using librosa library on python for our task. In our dataset, we have performances by various artists, each singing at different tonics. Tonic is just the reference note for construction of any Raga. Music performances in the same Raga can exhibit structural similarities when they use the same tonic. We performed tonic normalization on the extracted chromagram features, which aligns all songs to a common reference note, facilitating the classification model’s ability to discern Raga patterns irrespective of the original tonic used by the artists.
IV-B Tonic Normalisation
For a given audio signal x, let the short time fourier transform is given by $X(n,k)$. The center frequency of a pitch in an equal-tempered midi scale is given by:
$F_{pitch}(p)=440\times 2^{(p-69)/12}$ (1)
From these, we get a pitch-based log-frequency spectrogram, which converts the whole spectrogram having frequency axis equal to NFFT to have 128 values corresponding to the 128 midi frequencies. We basically assign each of the $X(n,k)$ to the closest $f_{coef(k)}$ frequency. So, for each of the 128 pitch values $p\in[0:127]$, we have:
$P(p)={k:F_{pitch}(p-0.5)<=F_{coef}(k)<F_{pitch}(p+0.5)}$ (2)
So, we have the log frequency spectrogram $Y_{LF}$ given by:
$Y_{LF}(n,p)=\sum_{k\in P(p)}|X(n,k)|^{2}$ (3)
To obtain chromagram features, we simply sum up the pitch values which differ by some integer multiple of an octave as they are perceived to be the same note musically:
$C(n,c)=\sum_{p\in[0:127]:pmod12=c}Y_{LF}(n,p)$ (4)
Now, to shift the chromagram representation of any song by a factor of "$a"$ semitones, we are essentially looking to get a new $X^{\prime}(n,k)$ such that:
$X^{\prime}(n,k)=X(n,k\times 2^{a/12})$ (5)
Shifting by "$a"$ semitones essentially means multiplying the frequency values in a spectrogram by $2^{a/12}$, where "$a"$ is a discrete value such that $a\in[1:11]$. This range for "$a"$ is chosen because a value greater than 11 would result in an octave shift, which is not a consideration for us since it would return to the same musical note. Now, we will get log-frequency spectrogram as:
$Y_{LF}^{\prime}(n,p)=\sum_{k\in P(p)}|X(n,k^{\prime})|^{2}$ (6)
where, $k^{\prime}=k\times 2^{a/12}$
Now, using Equation: 1,
$F_{pitch}(p)=k^{\prime}=440\times 2^{(p-69)/12}$ (7)
Substituting the value for $k^{\prime}$, and re-arranging terms, we get:
$k$ $=440\times 2^{((p-a)-69)/12}$ (8)
Comparing Equation: 8 with 7, we can see that for a given $k^{\prime}$, which gets shifted ${}^{\prime}a^{\prime}$ semitones, the set of frequency values given by $P(p^{\prime})$ is just all the frequency values present ${}^{\prime}a^{\prime}$ chromagram bins higher. i.e., the pitch values, represented by $F_{\text{pitch}}(p)$, undergo a shift when transitioning from $k^{\prime}$ to $k$ or vice versa, where the difference between the two pitches is given by the factor ${}^{\prime}a^{\prime}$. This shift signifies a displacement of the frequency values associated with the musical notes, particularly within the context of a chromagram bin. It implies that the set of pitch values $P(p)$, forming a bin in the chromagram, undergoes a uniform displacement by a factor of ${}^{\prime}a^{\prime}$. Consequently, while the bins themselves remain unchanged, the pitch values contained within each bin experience a consistent shift, maintaining the relative positioning of musical notes while adjusting their absolute frequencies. So, in practice, for applying tonic normalisation to our audio data, we just shift the extracted chromagram bins through the shorter dimension in a cyclic manner.
Figure 2: Model Architecture for CN2+LSTM+T
IV-C Classification Model Description
To determine the best performing model, we conduct an ablation study, experimenting with various CNN and LSTM-based architectures. Our first model, CN1+T, utilizes a simple CNN architecture with convolutional, max-pooling, and batch normalization layers, followed by a flatten layer, in a very similar way to a common Image classification model. Then the CN1+LSTM+T model extends this by replacing the flatten layer with an LSTM layer, ideal for processing time-series data like audio signals. Also, we provide tonic normalization as discussed in Section:IV-B and test the model performance. The CN2+LSTM model features a more complex CNN-LSTM architecture with 1D max-pooling to preserve temporal information, as shown in Fig: 2. Then the CN2+LSTM+T model just includes tonic normalization in the input features, following the exact configuration of CN2+LSTM model. The selection of layer configurations and activation functions is carefully optimized for classification performance. These models will be further detailed in sectionV, where we evaluate their effectiveness in ARI.
IV-D Classification Model Training
We train our DL models for the 12-class classification task, while using softmax activation in the last layer. Let $z_{i}$ represents the output logit for the $i^{th}$ neuron in the last dense layer having 12 neurons. Softmax function applied to these logits converts them into probability values as:
$p_{i}=softmax(z_{i})=\frac{e^{z_{i}}}{\sum_{j=1}^{N}e^{z_{j}}}$ (9)
Here, $p_{i}$ is the probability score for $i^{th}$ neuron in the layer, $N=12$ as we have a 12-class classification problem. Also, $i\in[1:N]$, so we get 12 probability values, such that $\sum_{i}{p_{i}}=1$. Then for training, we use categorical cross-entropy loss function given by:
$L_{CCE}=\sum_{i=1}^{N}t_{i}log(p_{i})$ (10)
Here, $t_{i}$ is the $i^{th}$ true label. In this case, $t_{i}$ will be 1 for only 1 neuron and 0 for all other neurons corresponding to a particular input. The training is carried out for multiple epochs until the model performances converges measured by precision, recall and f-1 score. In our experiments, we use cross-validation to assess the robustness and generalization ability of our models for the classification task.
IV-E Grad-CAM++
GradCAM++ builds upon GradCAM (Gradient-weighted Class Activation Maps), with a key advancement to it. Initially, we reload the CN2+LSTM+T model trained in above steps where the last CNN layer comprises 256 learned feature maps in our case as visible in the Fig: 2. To create a class-specific GradCAM++ activation map, each of these feature maps is weighted and then a weighted sum is taken to generate the final activation map or saliency map for the target class. We compute the weights of feature maps using the weights of those last CNN blocks, given by the equation:
$\alpha_{c}^{k}=\frac{1}{Z}\sum_{i}\sum_{j}\frac{\partial y^{c}}{\partial A^{k}_{ij}}$ (11)
Here, $A^{k}_{ij}$ represents each feature map (so we have $k$ feature maps and $k\in[1:256]$). Also, we have $c$ classes with $c\in[1:12]$. $\alpha_{c}^{k}$ represents the weights of $k^{th}$feature map for $c^{th}$ class. $Z$ is the dimension of each feature map, and $y^{c}$ represents the output logits whose gradients we compute using the chain rule. Using these computed weights as the weights of each of the $k$ feature maps, we get the final Grad-CAM attribution map, given by:
$L^{c}_{Grad-CAM}=ReLU\Bigl{(}\sum_{k}\alpha_{k}^{c}A^{k}\Bigl{)}$ (12)
$L^{c}_{Grad-CAM}$ represents the final attribution map for GradCAM, which is obtained by summing all the feature maps up-sampled to the original input size, given by $A^{k}$, with each $A^{k}$ weighted with a corresponding weight $\alpha_{k}^{c}$ according to a given output prediction class $c$. After that, ReLU is applied to keep only the positively impacting salience regions and remove the regions with negative gradient in the output saliency map or attribution map. But, to obtain the activation map for GradCAM++, we weight each pixel separately rather than weighting the whole feature map with a single weight. The new weights are given by:
$\alpha_{c}^{k}=\sum_{i}\sum_{j}w^{kc}_{ij}ReLU\Bigl{(}\frac{\partial y^{c}}{\partial A^{k}_{ij}}\Bigl{)}$ (13)
Here, $A^{k}_{ij}$ represents each pixel in the $k^{th}$ feature map. This $w^{kc}_{ij}$ aims to weigh each pixel separately, rather than the whole activation map. So, this $w^{kc}_{ij}$ is given by:
$w^{kc}_{ij}=\frac{\frac{\partial y^{c}}{(\partial A^{k}_{ij})^{2}}}{2\frac{\partial y^{c}}{(\partial A^{k}_{ij})^{2}}+\sum_{a}\sum_{b}A^{k}_{ab}\Bigl{(}\frac{\partial^{3}Y^{c}}{(\partial A^{k}_{ij})^{3}}\Bigl{)}}$ (14)
Using these, we finally obtain the attribution map for Grad-CAM++ given by:
$L^{c}_{Grad-CAM++}=ReLU\Bigl{(}\sum_{k}\alpha_{k}^{c}A^{k}\Bigl{)}$ (15)
This activation map given by $L^{c}_{Grad-CAM++}$ represents the desired visualisations for the regions of interest in a particular test example for a given class for the classifier model.
IV-F SoundLIME
SoundLIME[36] is specifically designed for audio, analogous to LIME’s[33] application in images. Here also, we need the pre-trained classifier model, but we treat it as a black-box model with no access to its gradients. Here, we temporally divide the chromagram into 30 equal parts or super-pixels, with one super-pixel representing the chromagram representation for each second of the audio. Next, we create 150 random perturbations for the given test input by applying random multi-hot masks of shape 30x1, meaning that each element of the mask either suppresses or retains each super-pixel. An exponential kernel is defined on L2 distance function $D(x,z)$ between the original image $x$ and the perturbed image $z$, denoted by $\pi_{x}(z)$:
$\pi_{x}(z)=exp(-D(x,z)^{2}/\sigma^{2})$ (16)
We train a weighted linear regression model, using the mask as input and querying the original classifier model for prediction of the masked input image of the chromagram, using that as the target.The significance of this approach can be understood from the fact that if most of the salient regions are already present in the masked input example and only non-salient regions are masked, the classifier model predicts the correct class. That becomes the label for our linear regression model that whether the given perturbed example results in the desired prediction or not. So, our weighted Linear regression model trained on these 150 perturbed data points learns the coefficients for each dimension of the mask and ultimately outputs which regions (super-pixels) of the mask are more salient for the prediction of the current test example.
Let $g\in G$ represents a family of interpretable models, in our case, $g$ represents the linear regression model. $f$ is the classifier model. $\omega(g)$ represents the complexity of the model, in our case it could be represented by the number of non-zero weights. The LIME explanation can be shown mathematically as:
$\zeta(x)=argmin_{g\in G}L(f,g,\pi_{x})+\omega(g)$ (17)
IV-G Saliency Evaluation Method
To evaluate the salient regions identified by the two ExAI models, we use the output of ExAI models, and compare them with the expert-annotated salient regions. We compute the precision for top 1 to 10 seconds for both the ExAI models. In our task, the concern lies solely with determining whether the top predicted regions are correct. Hence, precision is the best metric to employ in this situation, whereas recall and accuracy are not helpful. We also explain one correct and one incorrect prediction of the classifier and analyse their salient region. Also, the evaluation of correctness of ExAI model predictions is not required in our case, as we are solely focused on the salient regions’ correctness within the top few seconds in context of HCM, rather than assessing the ExAI model’s performance.
V Experiments and Results
V-A Data Pre-processing
To prepare the audio files for analysis, all non-musical speech has been removed. So, we firstly distribute the whole songs to train, test and validation sets, making sure that there is at least 1 song of each class in the test and validation sets. Then we split the songs into 30 second snippets for all the 3 splits with percentages of approximately 79%, 15%, and 6%, for training, validation, and test sets respectively as shown in Fig: 3 Notice that we split the dataset into train,test and validation sets first on file level, and then extract split them into 30-second clips. We do not directly create 30-seconds clips and then shuffle them as we have a possibility that we have snippets of the same song in train, test and validation set, which will add an additional bias in our dataset and will hinder the generalizability of our model.
Figure 3: Dataset pre-processing for the network
Also, the last part of each music recording is removed as it is mostly found to be having only 3 to 10 seconds of music, which may not even contain any salient region for getting classified to a particular raga and just serves the purpose of confusing the model. Also, generally in HCM performances, the endings of clips lasting 20 seconds or more may simply consist of a single note prolonged for a few seconds, which may not accurately represent the current Raga class. Thus, removing these endings should not adversely affect the analysis.
For our work, we extract chromagram features for these audio snippets, making each input sample of shape (938,12), having 938 time frames having 12 chromagram bins each.
All the songs, marked with their respective tonic by the music experts go through tonic normalisation as explained in IV-B. We make Note ”A” as the reference note for all and normalise accordingly. To achieve this, we just cyclically shift the obtained chromagram features through their shorter dimension throughout the 938 frames for each 30-sec file to obtain the desired normalisation.
V-B Classification results
We chose audio files belonging to 12 classes having the most number of audio files out of the dataset and train a 12-class classifier using the models discussed in IV-C. Classification results for our Ablation studies are shown in table: II (all values rounded off to 2 decimal places)
TABLE II: Combined Classification Report for Different Models
Classification Report for Different Models
Raga CN1+T CN1+LSTM+T CN2+LSTM CN2+LSTM+T
Precision Recall F1 Precision Recall F1 Precision Recall F1 Precision Recall F1
Bhairavi 0.89 0.53 0.67 0.83 0.74 0.78 0.0 0.0 0.0 0.93 0.91 0.92
Bihag 0.42 0.31 0.36 0.87 0.77 0.81 0.40 0.74 0.52 0.90 0.91 0.90
Des 0.09 0.08 0.08 0.63 0.74 0.68 0.06 0.08 0.07 0.78 0.66 0.70
Jog 0.85 0.95 0.90 0.97 0.86 0.91 0.88 0.88 0.88 0.91 0.89 0.90
Kedar 0.72 0.81 0.76 0.92 0.75 0.82 0.0 0.0 0.0 0.93 0.89 0.91
Khamaj 0.24 0.38 0.29 0.70 0.85 0.77 0.0 0.0 0.0 0.55 0.68 0.59
Malkauns 0.94 0.89 0.92 0.90 0.96 0.93 0.64 0.61 0.62 0.99 0.96 0.97
Maru-Bihag 0.57 0.56 0.56 0.80 0.90 0.84 0.52 0.68 0.60 0.88 0.85 0.86
Nayaki-Kanada 0.64 1.0 0.78 0.91 0.96 0.93 0.17 0.20 0.19 0.93 0.92 0.92
Shudha-Kalyan 0.95 0.37 0.53 0.79 0.75 0.77 0.67 0.67 0.67 0.89 0.94 0.92
Sohni 0.35 0.61 0.45 0.88 0.90 0.88 0.0 0.0 0.0 0.89 0.83 0.85
Yaman 0.66 0.86 0.75 0.87 0.81 0.84 0.68 0.52 0.60 0.90 0.91 0.90
Weighted Avg 0.68 0.64 0.63 0.85 0.83 0.83 0.50 0.50 0.50 0.90 0.89 0.89
V-B1 CN1+T
This model achieved an overall weighted f1-score of 0.78 on the test data for all 12 classes after cross-validation. The maximum and minimum f1-scores during cross-validation were 0.80 and 0.73, respectively. Clearly, this architecture is not be suitable for audio data, as flattening does not appropriately handle the dependent framewise information required for Raga detection in case of such time-series data.
V-B2 CN1+LSTM+T
This model showed improved classification performance with an f1-score of 0.83 after cross-validation. The maximum and minimum f1-score observed during cross-validation were 0.87 and 0.79, respectively. This model outperformed the CN1+T model, highlighting the importance of LSTM in stead of flattening in such scenarios.
V-B3 CN2+LSTM
Despite multiple training attempts and hyper-parameter tunings, this model did not perform well on the training data itself, even with a similar architecture to our main model (CNT+LSTM+T) or a even deeper CNN-LSTM model. This underscores the importance of tonic normalization for the ARI task. We also experimented with removing classes with very low f1-scores, simplifying the problem. While this led to some improvement in training, it suggests a lack of generalization to any number of Raga classes using this approach.
V-B4 CN2+LSTM+T
This model consistently achieved the highest performance among all models, with an f1-score of 0.89 after 6-fold cross-validation. For four classes, the CN1+LSTM+T model showed better results, however, the proposed model closely follows, demonstrating strong training across all 12 classes and better generalizability. The maximum and minimum f1-scores observed during cross-validation are 0.94 and 0.84, respectively.
In HCM performances involving Khayal or Dhrupad singing, the starting 5 to 10 minutes of a performance generally involves the ”Vilambhit” (very slow tempo) section, where the singer will be singing just straight notes stretched to a few seconds, until they start singing in ”Madhya Laya” (medium tempo of the range 100 to 150). Considering that this dataset is composed of actual HCM performances, it is expected that some of the 30 second chunks from the starting portions of such singing performances will have just a few straight notes stretched for longer periods of times. These kind of audio files (30-second) will most probably confuse the model since these would come out to be very similar for any ragas which have similar ”swaras” in their ”Aaroh” and ”Avroh”. Considering that, the achieved performance is remarkable. For our current analysis, we are focused on identifying salient regions based on predictions, so we report ARI results at the chunk-level only and do not predict Raga on song-level for our dataset.
Now, we carry out further analysis in terms of ExAI analysis to understand the credibility of this classifier.
Figure 4: Raga class distribution for test files
V-C Saliency Evaluation
V-C1 Saliency annotation
The music expert is given a randomly selected 64 audio files for which the model predicts the correct class. The task is to look for such regions in the audio file which would allow any hindustani music expert to judge if that audio snippet belongs to a particular Raga class. The Expert is asked to keep only those audio files which have at least 5 to 7 seconds of such region and at max 15 to 17 seconds. As it would make no sense to mark just one or two seconds, or the whole audio as the salient region. The salient region can be a single region, or may be 2 or 3 disjoint regions throught the 30 second audio chunk. Also, it was made sure that we have at least 3 audio examples per class. So after the annotation, we get 47 labelled test examples to evaluate our saliency methods. The distribution of test classes is shown in figure 4.
V-C2 Saliency evaluation
To evaluate the salient regions identified by the two ExAI models, we compare the predictions with the human-annotated saliency regions. For GC, we sum the activation map values across the frequency to get a 938x1 vector, showing saliency values for all 938 frames. We also convert the annotations of music experts into a multi-hot vector of length 938, where we mark 1 for salient region and 0 for non-salient. Now, to compare, we try out different windows starting from top 1 seconds upto top 10 seconds of GC prediction values and compute precision in each case.
For SL model, we get coefficients of linear regression model trained on the test data, in terms of 30 masks. We upscale this to get a multi-hot vector of 938x1 to match the dimensions of expert annotations. Now we compare the salient region given by this model in the same way as GC and compute precision values.
In both these cases, we compute the model’s precision for upto top 10 seconds out of the salient region, starting from 1 second. This is done so that we are able to judge how correct the top salient region of our model is. If the model is giving out decent results, it can be used as snippet for being representative of the particular Raga to use that for music teaching other applications.
TABLE III: Saliency report for GC and SL
Time frame GC Precision SL Precision
1 second 0.61 0.68
2 seconds 0.60 0.66
3 seconds 0.59 0.64
4 seconds 0.58 0.60
5 seconds 0.57 0.58
6 seconds 0.57 0.58
7 seconds 0.56 0.58
8 seconds 0.56 0.56
9 seconds 0.55 0.55
10 seconds 0.55 0.54
V-C3 Saliency results
Table: III shows the achieved average precision of explanations over the 47 test samples, compared with expert annotations for the most-salient N-seconds for both the ExAI models, where $N\in[1:10]$. Observe that for a smaller N, precision is high and it decreases as N increases, indicating that the top N seconds of highlighted salient regions by these ExAI models are very close to actual salient regions for that particular audio chunk, implying that the model is indeed relying on the time frames (the music in those time-frames) to make its predictions for current Raga class. Also, we can observe that SL shows higher performance which is also expected since GC considers activations only upto the CNN layers, leaving out the LSTM layers. Although there is no direct comparison metric among the two methods, we argue here that SoundLIME explanations will be more reliable as it takes predictions from the whole model (including the LSTM layers) while generating the explanations, hence these should prove to be a better representative of the predictions of the classification model.
TABLE IV: Analysis of individual samples
True Predicted softmax GC SL
Sample Sample probability Precision Precision
Des Des 0.95 0.77 0.8
Bihag Maru-Bihag 0.71 0.12 0.0
Figure 5: Salient Regions for Raga Des sample
Now, we study two individual test audios for salient regions. We select a ramdom correctly predicted test sample belonging to Raga class “Des”. The classifier shows a softmax probability score of 0.95 as shown in Table: IV with signifying that the model is quite confident about this prediction. The true salient region marked by the music expert, the salient region as reported by GC and SL are shown in Figures 5. The marked true region by the music expert came out to be around 15 seconds. We are concerned about finding how much of the region predicted by the model falls in the correct category. The precision score for GC and SL come out to be 0.77 and 0.8 respectively for 10 seconds, signifying a significant overlap and hence a significant amount of the prediction being based on the true concepts related to music only.
Figure 6: Salient Regions for Raga Bihag sample
We also conduct an individual study on two closely related Ragas in HCM, namely Bihag and Maru-Bihag. Maru-Bihag, being a musical blend of two Ragas, ”Maru” and ”Bihag”, may exhibit some characteristics of Raga Bihag in its composition, and vice versa. We can observe in the confusion matrix shown in Fig: 7 also that there are a few number of samples of class Bihag which have been mis-classified as Maru-Bihag.
Figure 7: Confusion Matrix for 12-class classifier
We carefully select such a false negative example and get it annotated for salient region showing the current Raga (Bihag in this case). Fig: 6 shows the true salient region, predicted salient region from GC and SL. We obtained top k seconds of salient region to match the length of salient region obtained from expert annotations from GC and SL models, shown in Fig: 6. The softmax probabilities as shown in table:IV of this model are 0.26 (true class: Bihag) and 0.71 (predicted class: Maru-Bihag), signifying that the model does recognise the true class (Bihag) to some extent in the given audio, but prediction of Maru-Bihag suggests that there is some ambiguous region in that particular audio snippet which confuses the model to differentiate between the two Raga classes. The same is also evident from the fact that the expert was only able to find a very small region from the whole audio that prominently shows the characteristics of Raga Bihag, implying that the audio chunk itself could be a bit ambiguous. Also, The precision score for GC and SL come out to be 0.12 and 0.0 respectively. We can clearly see in the figure6 that the salient regions of both GC and SL do not overlap the true region and points to totally different region, implying that there were frames which are similar to Raga Maru-Bihag in that audio sample and the frames implying Raga Bihag might not be that prominent, which can be confirmed by listening to the audio also.
VI Conclusion and Future Scope
In this paper, we introduce a novel labeled dataset for Hindustani Classical Music (HCM) analysis and achieve an f1-score of 0.89 for 12-class Raga classification task by carrying out Ablation studies. We observe that, providing tonic normalization to the extracted Chromagram features significantly enhances the performance of the Raga classifier. Following that, we employ two Explainable AI (ExAI) models to identify the salient regions for the classifier’s predictions and analyze their overlap with human annotations. This comparison reveals a significant overlap, affirming that the patterns or characteristics theoretically associated with the identification of a Raga for any audio are present and are being used by the Raga Identification model for the task. This serves as a correctness check for the music theory that specific motifs, phrases or characteristics are indeed responsible for the identification of any Raga by the Deep Learning model built on a dataset that represents actual HCM performances.
This work provides a valuable contribution to the field of explainability in Raga identification in Indian Art Music, setting the foundation for future research. The meticulously labeled dataset, including Ragas and tonic information, expands the landscape of available high-quality datasets in HCM analysis, and opens avenues for further research in music classification and analysis. Moving forward, Incorporating musicians’ perspectives and musical concepts like ”Khatka” and ”murki” to refine Raga classifiers can ensure correctness of Raga identification according to musical conventions and nuances.
Future research can focus on developing methods to generate listenable or musical explanations, which could serve as standalone identifiers of the Raga. This would include extracting phrases from audio clips that correspond to the Raga, offering potential applications in music pedagogy and beyond. It is important to note that experts may not always agree on the same set of phrases or regions, indicating the subjective nature of some aspects of music analysis. Another direction for future research can be to extend this work to other musical traditions and explore the generalizability of the approach.